OpenStudy (anonymous):

Solve the following word problem by using an equation and factoring. Write the solution only in the blank. The sum of a positive number and its square is 20. What is the number?

OpenStudy (anonymous):

i would guess and check on this one \[3^2+3=12\] too small try 4

OpenStudy (anonymous):

4 is too big .

OpenStudy (alfie):

it's -5

OpenStudy (anonymous):

actually i think 4 is just rigth

OpenStudy (alfie):

upps.. positive, fail

OpenStudy (anonymous):

\[4+4^2=4+16=20\]

OpenStudy (alfie):

4 is right ! :) a^2 + a -20 = 0. Solve for a, you get 4 and -5. You get the positive one!